Jon is considering the purchase of 340 acres of land for $1527.18 per acre. A bank will loan him $1,290 per acre of land and the loan will be fully amortized over 20 years at 18%. The outstanding balance of the loan will be paid at the end of the third year. Assume that the marginal tax rate is 20%. Suppose that the following information is given:

            Net Cash Flow before debt in Year 1= $162.00

            Net Cash Flow before debt in Year 2= $177.00

            Net Cash Flow before debt in Year 3= $1,522.00

(i) How much down payment is needed for the 340 acres?

(ii) Calculate the amount of loan payment per acre per period.

(iii) Calculate the per acre tax savings from paying interest in year 2.

(iv) How much money needs to be generated from other parts of the business to cover the cash deficit for the 340 acres in the 1st year, so that the investment is financially feasible?

(v) How much money needs to be generated from other parts of the business to cover the cash deficit for the 340 acres in the 2nd year, so that the investment is financially feasible?
Answer #1

Let period be 1 year i.e. yearly loan payments

Loan Amount = $1290*340 = $438600

Payment per year P can be calculated as

P/0.18*(1-1/1.18^20) = 438600

=> P =$81939.24 which is the loan payment per year

i) Down payment = 340* (1527.18-1290) = $80641.20

ii) Loan Payment per acre per year = $81939.24/340 =$240.9977 or $241

iii) Balance amount remaining at the end of 1 year = 438600*1.18- 81939.24 = $435608.76

So, Interest paid in 2nd year = 435608.76 * 0.18 = $78409.58

So, tax savings = 78409.58*0.2 = 15681.92

Tax Savings per acre = 15681.92/340 = $46.12

iv) Tax savings in 1st year = Interest paid * tax rate = ($438600* 0.18)*0.2 =$15789

Net cash flows generated in year 1= $162*340 = $55080

Total Cash flows = 55080+15789.60 = $70869.60

Requirement = $81939.24

So, Additional money required in year 1 = 81939.24 - 70869.60 = $11069.64

v) Tax savings in 2nd year = $15681.92

Net cash flows generated in year 2= $177*340 = $60180

Total Cash flows = 60180+15681.92 = $75861.92

Requirement = $81939.24

So, Additional money required in year 2 = 81939.24 - 75861.92 = $6077.32